Find the center (h,k) of the ellipse with the equation
9x^2-18x+4y^2+16y=11
Thanks:)

Answer by fractalier(6550)

Okay from
9x^2 - 18x + 4y^2 + 16y = 11
9(x^2 - 2x + 1) + 4(y^2 + 4y + 4) = 11 + 9 + 16
9(x - 1)^2 + 4(y + 2)^2 = 36
(x - 1)^2 / 4 + (y + 2)^2 / 9 = 1
center is at (1, -2)